OpenStudy (lexy148):

How to solve for t in d=16t^2

OpenStudy (anonymous):

\[d=16t^{2}\] \[\frac{ d }{ 16 }=t^{2}\] \[\sqrt{\frac{ d }{ 16 }}=\frac{ \sqrt{d} }{ 4 }\]

OpenStudy (lexy148):

i dont understand the whole sq root thing you did above @charliealbert

OpenStudy (anonymous):

chaeliealbert is correct, sorry about messing up above, swapped order of simplification

OpenStudy (anonymous):

Ohh \[\frac{ \sqrt{d} }{ 4 }=t\]

OpenStudy (anonymous):

I used the square root in both sides to make t to the first power

OpenStudy (lexy148):

oh okay thx :) @charliealbert

OpenStudy (anonymous):

No problem